Two planes are 1380 miles apart and traveling toward each other. One plane is traveling 80 mph faster than the other plane. The planes meet in 1.5 hours. Find the speed of each plane. Please show work.
Found 2 solutions by jorel1380, mananth:
Answer by jorel1380(3719)

1.5n+1.5(n+80)=1380
3n+120=1380
3n=1260
n=420
One plane is going 420 mph, the other is going 500 mph..

Answer by mananth(16946)

let one plane speed = x
other plane speed = x+80
combined speed = 2x+80
Distance = 1380
time = 1.5
D=rt
1380=(2x+80)1.5
1380/1.5= 2x+80
920=2x+80
2x=920-80
2x=840
x=420 mph speed of one plane
other plane 420+80 = 500 mph
m.ananth@hotmail.ca